OpenStudy (anonymous):

So i'm trying to study for a final exam. I found this practice problem online that said: What is the slope of a line parallel to the graph of 3x-5y=18. Am i suppose to use a slope intercept formula or something?

OpenStudy (anonymous):

frist write sope intercept from, u know how to writer it?

OpenStudy (anonymous):

so you do use it? or are there any other formulas that are also acceptable

OpenStudy (anonymous):

what formulas? I don't get the question u ask

OpenStudy (anonymous):

well, i have no idea how to do the problem. thats all, so i'm just guessing

OpenStudy (anonymous):

find slope or write slope intercept from pass 2 point or paralell form this equation 3x-5y=18.

OpenStudy (anonymous):

y=mx+b? y=3x+18? is that it

OpenStudy (anonymous):

y=mx+b? y=3x+18 is call slope intercept

OpenStudy (anonymous):

yeah okay. just asking if i plugged it in right.

OpenStudy (anonymous):

yes you r

OpenStudy (anonymous):

slpoe =3x if another line the same 3x than they are paralell

OpenStudy (anonymous):

yes you r soppese to

OpenStudy (anonymous):

\[\sqrt{?}\]vvvvvv

OpenStudy (anonymous):

okay so next step.is must be 3 over something.....

OpenStudy (anonymous):

3x-5y=18 y= (-3x+18)/-5 y = (3/5x)-(18/5)

OpenStudy (anonymous):

\[y=\frac{3x}{5}-\frac{18}{5}\]

OpenStudy (anonymous):

ok thanks! i see now

OpenStudy (anonymous):

u r welcom If I deserve medal then give me one.